OpenStudy (anonymous):

Hard conics question, need help ASAP! P, Q are two variable points on the hyperbola xy = c², such that the tangent at Q passes through A, the foot of the ordinate of P. (i.e. A is the intersection o the y-axis and the horizontal line through P.) Show that the locus of T, the intersection of tangents at P and Q, is a hyperbola with the same asymptotes as the given hyperbola. So confused... Please help?

OpenStudy (mathteacher1729):

Pretty sure \( xy = c^2\) is not an equation of a hyperbola. A hyperbola is of the form \[\frac{(x-h)^2}{a^2}\pm\frac{(y-k)^2}{b^2}=1\] Where \((h,k)\) is the center and \(a,b\) represent the lengths of the major & minor axes. (whichever is bigger is the "major").

OpenStudy (anonymous):

xy = c² is a rectangular hyperbola with branches in the first and third quadrants http://mathworld.wolfram.com/images/eps-gif/RectangularHyperbola_700.gif